# LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference

### v3

@@ 1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities Salton (1975). The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval Karpukhin et al. (2020) compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval Robertson et al. (1994) characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) Bai et al. (2020); Formal et al. (2021) and BM25 Robertson et al. (1994). Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores Wang et al. (2021) or ranks Cormack et al. (2009). Text retrieval is broadly applied in various scenarios, such as web search Lu et al. (2022), semantic textual similarity Bowman et al. (2015), fact checking Thorne et al. (2018), and retrieval augmented generation Lewis et al. (2020).
@@
Deploying retrieval systems requires different processing workflows for queries and documents. As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases Johnson et al. (2019); Robertson et al. (1994). Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere Yang et al. (2018a). On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Most existing works of text retrieval Karpukhin et al. (2020); Chen et al. (2024); Wang et al. (2024b) utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides. Recent state-of-the-art (SOTA) retrievers Wang et al. (2024b); BehnamGhader et al. (2024) employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources Kwon et al. (2023). Nowadays, LLM-based text retrieval incurs non-negligible online workloads, necessitating special optimizations for query inference efficiency.
Modern Transformers-based LMs handle text modeling with the attention mechanism Vaswani et al. (2017), including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts. As is the case for existing LLM-based text retrievers Wang et al. (2024b); BehnamGhader et al. (2024), queries and documents are modeled equally with the same LLM, enduring the same computational burdens. In contrast, traditional statistical retrieval algorithms, such as the popular BM25 Robertson et al. (1994), only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics. They do not handle the contextual traits and deep co-occurrences, and their encoders do not need to be deeply parametrized during inference. Experiments from Thakur et al. (2021) show that BM25 still performs well in many retrieval tasks.
Built on the above observations, to minimize query service cost, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture targeted at extreme lightweight query inference abilities. Our work retains a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities. But as is shown in Figure 1, we completely remove the deep modeling of queries. 1) For dense retrieval, our method end-to-end trains cacheable token vectors from the LLM-based query encoder. The dense query representations are simply aggregated by averaging corresponding token vectors shown in the texts. All query token vectors are pre-computed and cached into a simple Embedding layer. Query inference costs are drastically reduced to only one Embedding lookup. 2) For sparse retrieval, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency. is directly used as the sparse query vector. Thus, no more encoding stage for sparse retrieval is needed. The dense similarity scores are computed with cosine similarity, while the sparse scores directly use inner product similarity. These scores are summed Wang et al. (2021) for the final hybrid retrieval.
@@
LightRetriever is super fast and lightweight for query inference. Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR Thakur et al. (2021) and 8 Chinese tasks from CMTEB Retrieval Xiao et al. (2023)), maintaining an average of 95% retrieval performance.
@@ 2 Algorithm
Figure 2: Contrastive training of full-sized symmetric dense retrieval. A full-sized dual-encoder is used to model both queries and documents. Task-specific instructions are added as common practices Su et al. (2023); Wang et al. (2024b) to promote better domain adaptation abilities.
@@ 2.1 LM-based text retrieval
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ Chen et al. (2009). The above objective can be effectively optimized by the pairwise/listwise contrastive loss Cao et al. (2007), margin loss Rosset et al. (2003), etc. Following Karpukhin et al. (2020), listwise contrastive loss ($\ell^{CL}$) is used in our work,
@@
Similar to instructed fine-tuning (SFT) Ouyang et al. (2022), modern LM-based retrievers Su et al. (2023); Xiao et al. (2023); Wang et al. (2024a), especially LLM retrievers Wang et al. (2024b); BehnamGhader et al. (2024); Lee et al. (2024), customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although Su et al. (2023); Xiao et al. (2023) use instructions on both query and passage sides, most LLM-based retrievers Wang et al. (2024b); BehnamGhader et al. (2024); Lee et al. (2024) do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 Wang et al. (2024b) and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.6.
@@ 2.2 Dense retrieval of LightRetriever
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 Wang et al. (2024b). with the query encoder [^3] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.7.,
$v_{t_{i}}^{\text{den}}=Enc_{q}(Inst;t_{i})$ (3)
Then, the dense query vector $v_{q}^{\text{den}}$ of the whole query text is aggregated by averaging token vectors,
$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{\text{den}}$ (4)
@@ 2.3 Sparse retrieval of LightRetriever
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. Following SPLADE Formal et al. (2021), to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
@@
Here, $v_{d}^{\text{spr}}\in\mathbb{R}^{V}$ denotes the final sparse representation vector of the document. Given sparse vectors $v_{q}^{\text{spr}}$ and $v_{d}^{\text{spr}}$, the same contrastive loss in Equation 2 and auxiliary loss in Equation 9 are applied for representation learning.
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Following SPLADE Formal et al. (2021), given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator Paria et al. (2020) is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
Table 1: Comparison of retrieval speed and effectiveness. We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset Nguyen et al. (2016), evaluating both full symmetric retrievers and the query-lightweight LightRetriever. To assess the trade-off between efficiency and effectiveness, we also report nDCG@10 scores on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. For retrieval efficiency, we report the consumed time of: query tokenization, model encoding, maximum search time of Faiss/Lucene, and total end-to-end retrieval time. The overall throughput is measured in Queries Per Second (QPS).
@@ 3.1 Settings
Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity. To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO Ma et al. (2025) and Sentence Transformers Training Data Reimers (2019). English datasets include Amazon Review Ni et al. (2019), Eli5 Fan et al. (2019), FEVER Thorne et al. (2018), FiQA Maia et al. (2018), GooQA Khashabi et al. (2021), HotpotQA Yang et al. (2018b), MSMARCO Nguyen et al. (2016), NFCorpus Boteva et al. (2016), NPR Lucas et al. (2023), NQ Kwiatkowski et al. (2019), PAQ Lewis et al. (2021), Quora Duplicates Iyer et al. (2012), S2ORC Lo et al. (2020), Scifact Wadden et al. (2020), Specter Cohan et al. (2020), StackExchange Duplicates da Silva et al. (2018), E5 synthetic BeastyZ (2024), Trivia Joshi et al. (2017), WikiHow Koupaee and Wang (2018), and Yahoo Answers Zhang et al. (2015). Chinese datasets include cMedQA2 Zhang et al. (2018), DuReader Qiu et al. (2022), and T2Ranking Xie et al. (2023).
@@
BeIR Thakur et al. (2021) and CMTEB Retrieval Xiao et al. (2023) benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following Thakur et al. (2021); Xiao et al. (2023), accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
@@
Our experiments are conducted based on the implementation of tDRO Ma et al. (2025). All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE Formal et al. (2021), the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps Paria et al. (2020). LoRA Hu et al. (2022) is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B Touvron et al. (2023), and Qwen-2.5-1.5B, 3B, 7B Bai et al. (2023). The hybrid similarity scores are linearly summed from normalized dense and sparse scores Wang et al. (2021). All trainings are conducted on 8 NVIDIA H800 or A800 GPUs.
@@
To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCONguyen et al. (2016). All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search. Faiss and Anserini search in parallel, where Faiss searches more slowly in our cases.
@@ 3.2 Main results
Table 2 compares retrieval effectiveness across a wide range of retrievers on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding. Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers achieve the highest retrieval scores. For instance, Llama3.1-8b performs best on BeIR (56.8 nDCG@10), while Qwen2.5-7b reaches the best performance on CMTEB-R (70.1 nDCG@10). However, such performance comes at the cost of significant computational overhead, as discussed above.
@@
These results demonstrate that LightRetriever effectively balances retrieval quality and efficiency. By avoiding expensive query-side encoding and maintaining high retrieval performance, it establishes a promising new paradigm for scalable LLM-based retrieval systems, especially in latency-sensitive or high-throughput scenarios.

### v4

# LightRetriever: A LLM-based Text Retrieval Architecture with Extremely Faster Query Inference

@@ 1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [63]. The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval [30] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval [61] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 21] and BM25 [61]. Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [73] or ranks [15]. Text retrieval is broadly applied in various scenarios, such as web search [44], semantic textual similarity [8], fact checking [67], and retrieval augmented generation [40].
@@
Deploying retrieval systems requires different processing workflows for queries and documents. As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [28, 61]. Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [80]. On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Most existing works of text retrieval [30, 11, 72] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides. Recent state-of-the-art (SOTA) retrievers [72, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [38]. Despite the rapid progress on improving retrieval quality, little prior work has investigated the efficiency challenges of LLM-based retrievers, necessitating special optimizations for query inference efficiency.
Modern Transformer-based LMs empower text retrievers with rich contextual modeling through the attention mechanism [69], enabling deep interactions across contextual tokens. This capability explains the success of recent LLM-based dual-encoders [72, 5], but it also means that both queries and documents incur the same computational burden. In contrast, traditional statistical methods such as BM25 [61] rely purely on lexical overlap and token frequency, requiring negligible inference cost while still remaining competitive in many retrieval tasks [66]. This contrast highlights a practical tension: While document encoders can afford heavy pre-computation offline, query encoders must operate under strict online latency and resource constraints. Therefore, it is natural to ask: do queries truly need the same degree of deep contextual modeling as documents?
To strike a balance between performances and efficiencies, we propose LightRetriever, a new hybrid retrieval architecture that explicitly breaks the symmetry between document and query encoders. Our key insight is that while documents benefit from the full modeling power of LLMs, queries do not require equally heavy processing to achieve strong retrieval performance. To this end, LightRetriever preserves a full-sized LLM encoder on the document side, but entirely removes the deep modeling on the query side. 1) For dense retrieval, we end-to-end train cacheable token embeddings from the LLM-based query encoder and aggregate them through simple averaging, reducing query inference to a single embedding lookup; 2) For sparse retrieval, we directly map token counts to sparse vectors without any encoding. Dense and sparse scores are summed [73] for the final hybrid retrieval. LightRetriever achieves retrieval quality comparable to full dual-encoder LLMs, but with orders-of-magnitude faster query inference.
@@
LightRetriever is super fast and lightweight for query inference. Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [66] and 8 Chinese tasks from CMTEB Retrieval [77]), maintaining an average of 95% retrieval performance.
@@ 2 Algorithm
Figure 2: Contrastive training of full-sized symmetric dense retrieval. A full-sized dual-encoder is used to model both queries and documents. Task-specific instructions are added as common practices [65, 72] to promote better domain adaptation abilities.
@@ 2.1 LM-based text retrieval
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12]. The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [62], etc. Following [30], listwise contrastive loss ($\ell^{CL}$) is used in our work,
@@
Similar to instructed fine-tuning (SFT) [56], modern LM-based retrievers [65, 77, 71], especially LLM retrievers [72, 5, 39], customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although [65, 77] use instructions on both query and passage sides, most LLM-based retrievers [72, 5, 39] do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 [72] and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.6.
@@ 2.2 Dense retrieval of LightRetriever
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^1] Specifically, EOS pooling is used in our work by following Mistral-E5 [72]. with the query encoder [^2] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.7.,
$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{\text{den}}\text{, where }v_{t_{i}}^{\text{den}}=Enc_{q}(Inst;t_{i})$ (3)
@@ 2.3 Sparse retrieval of LightRetriever
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. To induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension by following [21],
@@
Here, $v_{d}^{\text{spr}}\in\mathbb{R}^{V}$ denotes the final sparse representation vector of the document. Given sparse vectors $v_{q}^{\text{spr}}$ and $v_{d}^{\text{spr}}$, the same contrastive loss in Equation 2 is applied for representation learning.
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [57] is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
Table 1: Comparison of retrieval speed and effectiveness. We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset [53], evaluating both full symmetric retrievers and the query-lightweight LightRetriever. To assess the trade-off between efficiency and effectiveness, we also report nDCG@10 scores on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. For retrieval efficiency, we report the consumed time of: query tokenization, model encoding, maximum search time of Faiss/Lucene, and total end-to-end retrieval time. The overall throughput is measured in Queries Per Second (QPS).
@@ 3.1 Settings
Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity. To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO [48] and Sentence Transformers Training Data [60]. English datasets include Amazon Review [54], Eli5 [20], FEVER [67], FiQA [49], GooQA [31], HotpotQA [81], MSMARCO [53], NFCorpus [7], NPR [45], NQ [37], PAQ [41], Quora Duplicates [26], S2ORC [43], Scifact [70], Specter [13], StackExchange Duplicates [17], E5 synthetic [4], Trivia [29], WikiHow [34], and Yahoo Answers [83]. Chinese datasets include cMedQA2 [82], DuReader [58], and T2Ranking [78].
@@
BeIR [66] and CMTEB Retrieval [77] benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following [66, 77], accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
@@
Our experiments are conducted based on the implementation of tDRO [48]. All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE [21], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps [57]. LoRA [25] is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B [68], and Qwen-2.5-1.5B, 3B, 7B [2]. The hybrid similarity scores are linearly summed from normalized dense and sparse scores [73]. All trainings are conducted on 8 NVIDIA H800 or A800 GPUs.
@@
To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCO[53]. All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search. Faiss and Anserini search in parallel, where Faiss searches more slowly in our cases.
@@ 3.2 Main results
Table 2 compares retrieval effectiveness on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding. Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers with deep LLM encoders achieve the best scores (e.g., Llama3.1-8b on BeIR with 56.8 nDCG@10, Qwen2.5-7b on CMTEB-R with 70.1 nDCG@10), but at significant computational cost.
LightRetriever achieves promising effectiveness while drastically reducing query-side computational burden. Across different backbones, it incurs only modest degradations (typically around 5% with 1–5 absolute points). For example, LightRetriever-Qwen2.5-7b achieves 53.8 nDCG@10 on BeIR and 66.5 nDCG@10 on CMTEB-R, only 2.8 and 3.6 lower than its full counterpart. Detailed task-level results are shown in Appendix Tables 911. Notably, LightRetriever retains strong performance on large-scale open-domain QA, fact-checking, and web search benchmarks such as FEVER [67], SciFact [70], TRECCOVID [16], and MSMARCO [53]. On smaller, domain-specific datasets (e.g., FiQA [49] and CmedQA [82]), drops can be moderately larger (around 5 points) but remain tolerable, suggesting good generalization with room for future task-specific adaptation.